OpenStudy (anonymous):

Which of the follow radian measures equals 747 degrees? A. 2π B. 2.08π C. 3.35π D. 4.15π E. 7.47π

OpenStudy (anonymous):

pi radian = 180 degree

OpenStudy (anonymous):

ok. how do i apply it to this?

OpenStudy (anonymous):

then, 1 degree=pi/180 radian so, 747 degree = pi/180 * 747 radian

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

solve for pi?

OpenStudy (anonymous):

nope, pi = 3.14

OpenStudy (anonymous):

use that to fin the value of 747 degree

OpenStudy (anonymous):

find

OpenStudy (anonymous):

srry.. didnt figure it out.

OpenStudy (anonymous):

747 degree = pi/180 * 747 radian =3.14/180 * 747 radian

OpenStudy (anonymous):

now can u?

OpenStudy (anonymous):

uh... i got 13.031

OpenStudy (anonymous):

so is it d?